OpenStudy (anonymous):

Compute the limit, which illustrates the use of l'Hopital's rule. lim x^x-x / 1-x + lnx x-->1

OpenStudy (zzr0ck3r):

parentheses please....

OpenStudy (zzr0ck3r):

\(\lim_{x\rightarrow 1}\frac{x^x-x}{1-x+\ln(x)}\)?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

you need then \(\lim_{x\rightarrow 1}\frac{\frac{d}{dx}(x^x-x)}{\frac{d}{dx}(1-x+\ln(x))}\) what is \(\frac{d}{dx}(x^x-x)\)?

OpenStudy (anonymous):

x^x (ln(x) +1) -1

OpenStudy (anonymous):

i had got the overall answer -1 but it was wrong

OpenStudy (zzr0ck3r):

and \(\frac{d}{dx}(1-x+\ln(x))\)?

OpenStudy (anonymous):

1/x -1

OpenStudy (zzr0ck3r):

so we have \(\lim_{x\rightarrow 1}\frac{x^x\ln(x)+x^x-1}{\frac{1-x}{x}}\) so again we get 0, so now we do it again \(\frac{d}{dx}(x^x\ln(x)+x^x-1)\)?

OpenStudy (zzr0ck3r):

so again we get 0/0 is what I meant...

OpenStudy (anonymous):

x^x(1/x(ln(x) +1^2)

OpenStudy (zzr0ck3r):

I get \(x^x(\frac{1}{x}+(\ln(x)+1)^2)\)

OpenStudy (anonymous):

ya sorry thats what i meant

OpenStudy (anonymous):

and for the other ones its -1/x^2

OpenStudy (zzr0ck3r):

plug in 1 to the top you get 2, so -2 is the answer

OpenStudy (anonymous):

okay thanks for some reason i had gotten -1 but i checked again now and i see that it is -2

OpenStudy (zzr0ck3r):

dont you hate when you do all the hard stuff and it was the easy part you messed up....story of my life

OpenStudy (anonymous):

lol agreed